# Variational Search Distributions

### v2

@@ Abstract
We develop variational search distributions (VSD), a method for finding discrete, combinatorial designs of a rare desired class in a batch sequential manner with a fixed experimental budget. We formalize the requirements and desiderata for this problem and formulate a solution via variational inference. In particular, VSD uses off-the-shelf gradient based optimization routines, can learn powerful generative models for designs, and can take advantage of scalable predictive models. We derive asymptotic convergence rates for learning the true conditional generative distribution of designs with certain configurations of our method. After illustrating the generative model on images, we empirically demonstrate that VSD can outperform existing baseline methods on a set of real sequence-design problems in various biological systems.
@@ 1 Introduction
We consider a variant of the active search problem (Garnett et al., 2012; Jiang et al., 2017; Vanchinathan et al., 2015), where we wish to find as many members (designs) of a rare desired class in a batch sequential manner with a fixed experimental budget. Examples of this are compounds that could be useful pharmaceutical drugs, or highly active enzymes for catalyzing chemical reactions. We assume the design space is discrete or partially discrete, high-dimensional, and practically innumerable. For example, the number possible configurations of a single protein is $20^{{\mathcal{{O}}}(100)}$ (see, e.g., Sarkisyan et al., 2016).
We are interested in this objective for a variety of reasons. We may wish to study the properties of the “fitness landscape” (Papkou et al., 2023) to gain a better scientific understanding of a phenomenon such as natural evolution. Or, we may not be able to completely specify the constraints and objectives of a task, but we would like to characterize the space of feasible designs. For example, we want enzymes that can degrade plastics in an industrial setting, but we may not yet know the exact conditions (e.g. temperature, pH), some of which may be anti-correlated with enzyme catalytic activity.
Assuming we can take advantage of a prior distribution over designs, we formulate the search problem as inferring the posterior distribution over rare, desirable designs. Importantly, this posterior can be used for generating new designs. Specifically, we use (black-box) variational inference (VI) (Ranganath et al., 2014), and so refer to our method as variational search distributions (VSD). Our major contributions are: (1) we formulate the batch active search objective over a (practically) innumerable discrete design space as an instance of variational inference, (2) we present an algorithm, VSD, which solves this objective, and (3) we show that VSD satisfies well-defined requirements and desiderata specific to our problem. For example, it uses off-the-shelf gradient based optimization routines, is able to learn powerful generative models, and can take advantage of scalable predictive models. In our experiments we show that VSD can outperform existing baseline methods on a set of real applications. Finally, we evaluate our approach on the related sequential black-box optimization (BBO) problem, where we want to find the globally optimal design for a specific objective and show competitive performance when compared with state-of-the-art methods, e.g., based on Bayesian optimization (BO) (Garnett, 2023).
@@ 2 Method
2.1 Problem Formulation
We are given a design space $\mathcal{X}$, which can be discrete or mixed discrete-continuous and high dimensional, and where for each instance that we choose $\mathbf{x}\in\mathcal{X}$, we measure some corresponding property of interest (so-called fitness) $y\in$. For example, in our motivating application of DNA/RNA or protein sequences (henceforth referred to as just sequences), $\mathcal{X}=\mathcal{V}^{M}$ where $\mathcal{V}$ is the sequence vocabulary (e.g., amino acid labels, $|\mathcal{V}|=20$) and $M$ is the length of the sequence. However, we do not limit the application of our method to sequences. Using this framing, a real world experiment (for example, measuring the activity of an enzyme) can be modeled as an unknown relationship,
@@ 2.1 Problem Formulation
for some black-box function (e.g. the experiment), ${f\!\centerdot}$, and measurement error ϵitalic-$\epsilon\in$, distributed according to $p({\epsilon})$ with $\mathbb{E}_{p({\epsilon})}\!\left[{\epsilon}\right]=0$. Instead of wanting to model the whole space, we are only interested in a set of events which we choose based on fitness $y$. For instance (refer to Figure 1), we may be interested in the fittest measurable design; all designs above a minimum level of feasibility (e.g. a wild-type sequence), $\tau\in$; the distribution of these feasible designs; or the shape of the black-box function for these feasible designs,
$\mathbf{x}^{*}=\operatorname*{arg\,\!max}_{\mathbf{x}}{f\!\centerdot}(\mathbf{x})+\epsilon,\quad\mathcal{S}:=\{\mathbf{x}:y>\tau\},\quad p({\mathbf{x}}|{y>\tau}),\quad\textrm{or}\quad\mathcal{F}:=\{{f\!\centerdot}(\mathbf{x}):\mathbf{x}\in\mathcal{S}\}.$ (2)
Our primary focus in this work is to estimate the super level-set distribution $p({\mathbf{x}}|{y>\tau})$ in a sequential manner. We assume that $\mathcal{S}$ are rare events in a high dimensional space, and that we have access to a prior belief, $p({\mathbf{x}})$, which helps narrow in on this subset of $\mathcal{X}$. We are given a dataset, $\mathcal{D}_{N}:=\{(y_{n},\mathbf{x}_{n})\}_{n=1}^{N}$, which may contain only a few instances of $y_{n}>\tau$. Given $p({\mathbf{x}})$ and $\mathcal{D}_{N}$ we aim to recommend batches of unique candidates, $\{\mathbf{x}_{bt}\}^{B}_{b=1}$, for experimental evaluation (Equation 1) in a series of rounds, $t\in\{1,\ldots,T\}$, where $B={\mathcal{{O}}}(1000)$ and we desire $\mathbf{x}_{bt}\in\mathcal{S}$. Each round, $\mathcal{D}_{N}$ is augmented with the experimental results of the previous batch, so $N\leftarrow N+B$. Estimating this super level-set distribution of $\mathbf{x}$ is computationally and statistically challenging and, therefore, we cast this as a variational inference problem. As we shall see later, our solution allows us to satisfy the following requirements and additional desiderata for our problem.
@@
Problem requirements (R) and other desiderata (D).
@@
feasible designs, $y>\tau$, are rare events in $\mathcal{X}$ that need to be identified
@@
non-myopic candidate generation, $\mathbf{x}\in\mathcal{S}\subset\mathcal{X}$, for sequential experiments
@@
(D3) Generative
models, $\mathbf{x}^{(s)}\sim q({\mathbf{x}})$, that are task-specific for fit designs
@@
predictive models that enable high-throughput experiments.
(a) $\operatorname*{arg\,\!max}_{\mathbf{x}}y(\mathbf{x})$
(b) $\mathcal{S}$
@@
(e) $\mathcal{F}$
Figure 1: Fitness landscape tasks. (1(a)) A noise-less fitness landscape, ${f\!\centerdot}(\mathbf{x})$ and ‘$\times$’ – the maximum fitness design, $\mathbf{x}^{*}$. (1(b)) The super level-set of all fit designs – white hatched area, $\mathcal{S}$. (1(c)) Prior belief $p({\mathbf{x}})$. (1(d)) The density/mass function of the super level-set, $p({\mathbf{x}}|{y>\tau})$ – blue contours. (1(e)) The black box function for the super level-set, $\mathcal{F}$. See Equation 2 for definitions of these these tasks. Our primary goal is to estimate the density or mass function of the super level-set, (1(d)). Since we assume a noisy relationship between ${f\!\centerdot}$ and $y$, the super level-set will not have a hard boundary as depicted, and $p({\mathbf{x}}|{y>\tau})$ will be non-zero over all $\mathcal{X}$.
@@ 2.2 Variational Search Distributions
where $q({\mathbf{x}}|{\phi})$ is a parameterized distribution from which we sample experimental candidate designs $\mathbf{x}_{bt}$, (D3) Generative, and which we aim to match to $p({\mathbf{x}}|{y>\tau})$. The difficulty is that we cannot directly evaluate or empirically sample from $p({\mathbf{x}}|{y>\tau})$. However, if we consider the reverse Kullback-Leibler (KL) divergence,
@@
where we have expanded $p({\mathbf{x}}|{y>\tau})$ using Bayes rule and dropped the constant term $p({y>\tau})$, we note that we no longer require evaluation of $p({\mathbf{x}}|{y>\tau})$ directly. We recognize the right hand side of Equation 4 as the well known (negative) variational evidence lower bound (ELBO),
@@
For this we assume access to a prior distribution over the space of designs, $p({\mathbf{x}})$, that may be informed from the data at hand. Henceforth, as we will develop a sequential algorithm, we will denote this prior with $p({\mathbf{x}}|{\mathcal{D}_{0}})$. We note the relationship between $\log p({y>\tau}|{\mathbf{x}})$ and the probability of improvement (PI) acquisition function from BO (Kushner, 1964),
$\log p({y>\tau}|{\mathbf{x}}):=\log\mathbb{E}_{p({y}|{\mathbf{x},\mathcal{D}_{N}})}\!\left[{\mathbbm{1}[{y>\tau}]}\right]=\log\alpha_{PI}\!\left({\mathbf{x},\mathcal{D}_{N},\tau}\right).$ (6)
@@
The method that maximizes the objective in Equation 7 we call variational search distributions (VSD), since we are using the variational posterior distribution as a means of searching the space of fit designs, satisfying (R1) Rare, (R2) Sequential and (R4) Batch. It is well known that when the true posterior is a member of the variational family indexed by $\phi$, the above variational inference procedure has the potential to recover the exact posterior distribution. To recommend experimental candidates we sample a set of designs from our search distribution each round,
@@
We discuss the relationship between VSD and BO in Appendix F. In general, because of the discrete combinatorial nature of our problem, we cannot use the re-parameterization trick (Kingma & Welling, 2014) to estimate the gradients of the ELBO. Instead, we use the score function gradient estimator (Williams, 1992; Mohamed et al., 2020) with standard gradient descent methods (D2) Gradient,
$\nabla_{\phi}\mathcal{L_{\textrm{ELBO}}}\!\left({\phi,\tau,\mathcal{D}_{N}}\right)$ $=\mathbb{E}_{q({\mathbf{x}}|{\phi})}\!\left[{\left(\log\alpha_{PI}\!\left({\mathbf{x},\mathcal{D}_{N},\tau}\right)-\log\frac{q({\mathbf{x}}|{\phi})}{p({\mathbf{x}}|{\mathcal{D}_{0}})}\right)\nabla_{\phi}\log q({\mathbf{x}}|{\phi})}\right],$ (9)
where we use Monte-Carlo sampling to approximate this expectation with a suitable variance reduction scheme, such as using a control variate or baseline (Mohamed et al., 2020). We find that the exponentially smoothed average of the ELBO works well in practice, and is the same strategy employed in Daulton et al. (2022). Effectively, VSD implements black-box variational inference (Ranganath et al., 2014) for parameter estimation, and despite the high-dimensional nature of $\mathcal{X}$, we find we only need ${\mathcal{{O}}}(1000)$ samples to estimate the required expectations for ELBO optimization on problems with $M=\mathcal{O}(100)$, satisfying (R3) Discrete. Note that Equation 79 do not involve any data ($\mathcal{D}_{N}$) directly, only indirectly through the acquisition function. Hence the scalability of VSD is dependent on the complexity of training the underlying estimator of $p({y}|{\mathbf{x},\mathcal{D}_{N}})$.
@@ 2.3 Class Probability Estimation
So far our method indirectly computes the PI by transforming the predictions of a GP surrogate model, $p({y}|{\mathbf{x},\mathcal{D}_{N}})$, as in Equation 6. Instead we may choose to follow the reasoning used by Bayesian optimization by density-ratio estimation (BORE) in Tiao et al. (2021); Oliveira et al. (2022); Song et al. (2022), and directly estimate the quantity we care about, $p({y>\tau}|{\mathbf{x},\mathcal{D}_{N}})$. We do this with class probability estimation (CPE) on the labels $z:=\mathbbm{1}[{y>\tau}]\in\{0,1\}$ so $p({y>\tau}|{\mathbf{x},\mathcal{D}_{N}})=p({z=1}|{\mathbf{x},\mathcal{D}_{N}})\approx\pi_{\theta}({\mathbf{x}})$, where $\pi_{\theta}:\mathcal{X}\to[0,1]$. We can recover the class probability estimates using a proper scoring rule (Gneiting & Raftery, 2007) such as Brier score or log-loss on training data, $\mathcal{D}_{N}^{z}=\{(z_{n},\mathbf{x}_{n})\}^{N}_{n=1}$, e.g.,
The VSD objective using CPE becomes,
@@
into which we plug $\theta_{t}^{*}=\operatorname*{arg\,\!min}_{\theta}\mathcal{L_{\textrm{CPE}}}\!\left({\theta,\mathcal{D}_{N}^{z}}\right)$. Using a CPE also opens up the choice of estimators that are more scalable than GP-PI, satisfying our desiderata (D4) Scalable. This may be crucial if we choose to run more than a few rounds of experiments with $B={\mathcal{{O}}}(1000)$. Additionally, since VSD is a black-box method, we can choose to use CPEs that are non-differentiable, such as decision tree ensembles. The complete VSD algorithm is given in Algorithm 1, in which we have allowed for a threshold function, $\tau_{t}=f_{\tau}\!\left({\{y:y\in\mathcal{D}_{N}\},\gamma_{t}}\right)$. This function can be used to modify the threshold each round, e.g. following Tiao et al. (2021), an empirical quantile function $\tau_{t}=\hat{Q}_{y}(\gamma_{t})$ where $\gamma_{t}\in(0,1)$, or a constant $\tau$ in the case of estimating the distribution of the super level-set.
Algorithm 1 VSD optimization loop with CPE.
1: Threshold $\gamma_{1}$ and $f_{\tau}$, dataset $\mathcal{D}_{N}$, black-box ${f\!\centerdot}$, prior $p({\mathbf{x}}|{\mathcal{D}_{0}})$, CPE $\pi_{\theta}({\mathbf{x}})$, variational family $q({\mathbf{x}}|{\phi})$, budget $T$ and $B$.
@@
3:     $\mathcal{D}^{z}_{N}$leftarrow $\{(z_{n},\mathbf{x}_{n})\}^{N}_{n=1}$,  where $z_{n}=\mathbbm{1}[{y_{n}>\tau}]$
@@
12:     $\mathcal{D}_{N+B}$leftarrow $\mathcal{D}_{N}\cup\{(\mathbf{x}_{bt},y_{bt})\}^{B}_{b=1}$

### v4

@@ Abstract
We develop variational search distributions (VSD), a method for conditioning a generative model of discrete, combinatorial designs on a rare desired class by efficiently evaluating a black-box (e.g. experiment, simulation) in a batch sequential manner. We call this task active generation; we formalize active generation’s requirements and desiderata, and formulate a solution via variational inference. VSD uses off-the-shelf gradient based optimization routines, can learn powerful generative models for desirable designs, and can take advantage of scalable predictive models. We derive asymptotic convergence rates for learning the true conditional generative distribution of designs with certain configurations of our method. After illustrating the generative model on images, we empirically demonstrate that VSD can outperform existing baseline methods on a set of real sequence-design problems in various protein and DNA/RNA engineering tasks.
@@ 1 Introduction
We consider a variant of the active search problem (Garnett et al., 2012; Jiang et al., 2017; Vanchinathan et al., 2015), where we wish to find members (designs) of a rare desired class in a batch sequential manner with a fixed black-box evaluation (e.g. experiment) budget. We call sequential active learning of a generative model of these designs active generation. Examples of rare designs are compounds that could be useful pharmaceutical drugs, or highly active enzymes for catalyzing chemical reactions. We assume the design space is discrete or partially discrete, high-dimensional and practically innumerable. For example, the number of possible configurations of a single protein is $20^{{\mathcal{{O}}}(100)}$ (see, e.g., Sarkisyan et al., 2016). Learning a generative model of these designs allows us to circumvent the need for traversing the whole search space.
We are interested in this active generation objective for a variety of reasons. We may wish to study the properties of the “fitness landscape” (Papkou et al., 2023) to gain a better scientific understanding of a phenomenon such as natural evolution. Or, we may not be able to completely specify the constraints and objectives of a task, but we would like to characterize the space of, and generate new feasible designs. For example, we want enzymes that can degrade plastics in an industrial setting, but we may not yet know the exact conditions (e.g. temperature, pH), some of which may be anti-correlated with enzyme catalytic activity. Alternatively, if we know these multiple objectives and constraints, we may only want to generate designs from a Pareto set.
Assuming we can take advantage of a prior distribution over designs, we formulate the search problem as inferring the posterior distribution over rare, desirable designs. Importantly, this posterior can be used for generating new designs. Specifically, we use (black-box) variational inference (VI) (Ranganath et al., 2014), and so refer to our method as variational search distributions (VSD). Our major contributions are: (1) we formulate the batch active generation objective over a (practically) innumerable discrete design space, (2) we present a variational inference algorithm, VSD, which solves this objective, (3) we show that VSD performs well theoretically and empirically, and (4) we connect active generation to other recent advances in black-box optimization (BBO) of discrete sequences that use generative models. VSD uses off-the-shelf gradient based optimization routines, is able to learn powerful generative models, and can take advantage of scalable predictive models. In our experiments we show that VSD can outperform existing baseline methods on a set of real applications. Finally, we evaluate our approach on the related sequential BBO problem, where we want to find the globally optimal design for a specific objective and show competitive performance when compared with state-of-the-art methods, e.g., based on latent space optimization (LSO) (Gruver et al., 2023).
@@ 2 Method
2.1 The Problem of Active Generation
We are given a design space $\mathcal{X}$, which can be discrete or mixed discrete-continuous and high dimensional, and where for each instance that we choose $\mathbf{x}\in\mathcal{X}$, we measure some corresponding property of interest (so-called fitness) $y\in$. For example, in our motivating application of DNA/RNA or protein sequences, $\mathcal{X}=\mathcal{V}^{M}$, where $\mathcal{V}$ is the sequence vocabulary (e.g., amino acid labels, $|\mathcal{V}|=20$) and $M$ is the length of the sequence. However, we do not limit the application of our method to sequences. Using this framing, a real world experiment (e.g., measuring the activity of an enzyme) can be modeled as an unknown relationship,
@@ 2.1 The Problem of Active Generation
for some black-box function (the experiment), ${f\!\centerdot}$, and measurement error ϵitalic-$\epsilon\in$, distributed according to $p({\epsilon})$ with $\mathbb{E}_{p({\epsilon})}\!\left[{\epsilon}\right]=0$. Instead of modeling the whole space, $\mathcal{X}$, we are only interested in a set of events which we choose based on fitness, $\mathcal{S}\subset\mathcal{X}$. In particular for active generation we wish to learn a generative model, $q({\mathbf{x}})$, that only returns samples $\mathbf{x}^{(s)}\in\mathcal{S}$ by efficiently querying the black-box function in Equation 1. We assume that $\mathcal{S}$ are rare events in a high dimensional space, and that we have access to a prior belief, $p({\mathbf{x}})$, which helps narrow in on this subset of $\mathcal{X}$. We are given an initial dataset, $\mathcal{D}_{N}:=\{(y_{n},\mathbf{x}_{n})\}_{n=1}^{N}$, which may contain only a few instances of $\mathbf{x}_{n}\in\mathcal{S}$. Given $p({\mathbf{x}})$ and $\mathcal{D}_{N}$ we aim to generate batches of unique candidates, $\{\mathbf{x}_{bt}\}^{B}_{b=1}$, for black-box (experimental) evaluation in a series of rounds, $t\in\{1,\ldots,T\}$, where $B={\mathcal{{O}}}(1000)$ and we desire $\mathbf{x}_{bt}\in\mathcal{S}$. After each round $\mathcal{D}_{N}$ is augmented with the black-box results of the batch, i.e. $\mathcal{D}_{N}\leftarrow\mathcal{D}_{N}\cup\{(\mathbf{x}_{bt},y_{bt})\}^{B}_{b=1}$. As we shall see later, our solution allows us to satisfy the following requirements and additional desiderata for active generation.
@@
Active generation requirements (R) and other desiderata (D).
@@
feasible designs, $\mathcal{S}$, are rare events in $\mathcal{X}$ that need to be identified
@@
non-myopic candidate generation, $\mathbf{x}\in\mathcal{S}$, for sequential black-box evaluation
@@
(R5) Generative
models, $\mathbf{x}^{(s)}\sim q({\mathbf{x}})$, that are task-specific for rare, feasible designs
@@
predictive models that enable high-throughput evaluation/experiments.
Like active search (Garnett et al., 2012) in our case we are interested in the solution space of the super level-set, $\mathcal{S}_{\textrm{SLS}}:=\{\mathbf{x}:{f\!\centerdot}(\mathbf{x})>\tau\}$ for a threshold $\tau\in$ (e.g., wild-type fitness). As we only have access to noisy measurements, $y$, our task is to estimate the super level-set distribution, $p({\mathbf{x}}|{y>\tau})$, using active generation. Estimating this distribution over $\mathcal{S}_{\textrm{SLS}}$ is computationally and statistically challenging and, therefore, we cast this as a variational inference problem. We also consider the case of BBO for which $\mathcal{S}_{\textrm{BBO}}:=\operatorname*{arg\,\!max}_{\mathbf{x}}{f\!\centerdot}(\mathbf{x})$, and we show that we can accommodate this in our variational framework by iteratively raising $\tau_{t}$ per round, $t$. We visualize the properties and models involved in active generation as applied to a continuous “fitness landscape” in Figure 1.
(a) $\operatorname*{arg\,\!max}_{\mathbf{x}}{f\!\centerdot}(\mathbf{x})$
(b) $\{\mathbf{x}:{f\!\centerdot}(\mathbf{x})>\tau\}$
@@
Figure 1: Fitness landscape properties and models. (1(a)) A noise-less fitness landscape, ${f\!\centerdot}(\mathbf{x})$, and the maximum fitness design, $\mathcal{S}_{\textrm{BBO}}=\{\mathbf{x}^{*}\}$, as the white ‘$\times$’. (1(b)) The super level-set, $\mathcal{S}_{\textrm{SLS}}$, of all fit designs as the white hatched area. (1(c)) Prior belief $p({\mathbf{x}})$. (1(d)) The density/mass function of the super level-set, $p({\mathbf{x}}|{y>\tau})$, as blue contours. Our goal is to sequentially estimate a generative model for the distribution of the super level-set (1(d)). We assume a noisy relationship between ${f\!\centerdot}$ and $y$, so the super level-set will not have a hard boundary, and $p({\mathbf{x}}|{y>\tau})$ will be defined over all $\mathcal{X}$.
@@ 2.2 Variational Search Distributions
where $q({\mathbf{x}}|{\phi})$ is a parameterized distribution from which we sample candidate designs $\mathbf{x}_{bt}$, (R5) Generative, and which we aim to match to $p({\mathbf{x}}|{y>\tau})$. The difficulty is that we cannot directly evaluate or empirically sample from $p({\mathbf{x}}|{y>\tau})$. However, if we consider the reverse Kullback-Leibler (KL) divergence,
@@
where we have expanded $p({\mathbf{x}}|{y>\tau})$ using Bayes rule and dropped the constant term $p({y>\tau})$, we note that we no longer require evaluation of $p({\mathbf{x}}|{y>\tau})$ directly. We recognize the right-hand side of Equation 3 as the well known (negative) variational evidence lower bound (ELBO),
@@
For this we assume access to a prior distribution over the space of designs, $p({\mathbf{x}})$, that may be informed from the data at hand (or pre-trained). Henceforth, as we will develop a sequential algorithm, we will denote this prior as $p({\mathbf{x}}|{\mathcal{D}_{0}})$. We note the relationship between $\log p({y>\tau}|{\mathbf{x}})$ and the probability of improvement (PI) acquisition function from Bayesian optimization (BO) (Kushner, 1964),
$\log p({y>\tau}|{\mathbf{x}})\approx\log p({y>\tau}|{\mathbf{x},\mathcal{D}_{N}})=\log\mathbb{E}_{p({y}|{\mathbf{x},\mathcal{D}_{N}})}\!\left[{\mathds{1}[{y>\tau}]}\right]=\log\alpha_{PI}\!\left({\mathbf{x},\mathcal{D}_{N},\tau}\right).$ (5)
@@
We refer to the method that maximizes the objective in Equation 6 as variational search distributions (VSD), since we are using the variational posterior distribution as a means of searching the space of fit designs, satisfying (R1) Rare, (R2) Sequential and (R4) Batch. It is well known that when the true posterior is a member of the variational family indexed by $\phi$, the above variational inference procedure has the potential to recover the exact posterior distribution. To recommend candidates for black-box evaluation we sample a set of designs from our search distribution each round,
@@
We discuss the relationship between VSD and BO in Appendix G. In general, because of the discrete combinatorial nature of our problem, we cannot use the re-parameterization trick (Kingma & Welling, 2014) to estimate the gradients of the ELBO straightforwardly. Instead, we use the score function gradient, also known as REINFORCE (Williams, 1992; Mohamed et al., 2020) with standard gradient descent methods (D2) Gradient such as Adam (Kingma & Ba, 2014),
$\nabla_{\phi}\mathcal{L_{\textrm{ELBO}}}\!\left({\phi,\tau,\mathcal{D}_{N}}\right)$ $=\mathbb{E}_{q({\mathbf{x}}|{\phi})}\!\left[{\left(\log\alpha_{PI}\!\left({\mathbf{x},\mathcal{D}_{N},\tau}\right)-\log q({\mathbf{x}}|{\phi})+\log p({\mathbf{x}}|{\mathcal{D}_{0}})\right)\nabla_{\phi}\log q({\mathbf{x}}|{\phi})}\right].$ (8)
Here we use Monte-Carlo sampling to approximate the expectation with a suitable variance reduction scheme, such as control variates (Mohamed et al., 2020). We find that the exponentially smoothed average of the ELBO works well in practice, and is the same strategy employed in Daulton et al. (2022). VSD implements black-box variational inference (Ranganath et al., 2014) for parameter estimation, and despite the high-dimensional nature of $\mathcal{X}$, we find we only need ${\mathcal{{O}}}(1000)$ samples to estimate the required expectations for ELBO optimization on problems with $M=\mathcal{O}(100)$, satisfying (R3) Discrete. Note that Equation 68 do not involve any data ($\mathcal{D}_{N}$) directly, only indirectly through the acquisition function. Hence the scalability of VSD is dependent on the complexity of training the underlying estimator of $p({y}|{\mathbf{x},\mathcal{D}_{N}})$.
@@ 2.3 Class Probability Estimation
So far our method indirectly computes PI by transforming the predictions of a GP surrogate model, $p({y}|{\mathbf{x},\mathcal{D}_{N}})$, as in Equation 5. Instead we may choose to follow the reasoning used by Bayesian optimization by density-ratio estimation (BORE) in Tiao et al. (2021); Oliveira et al. (2022); Song et al. (2022), and directly estimate the quantity we care about, $p({y>\tau}|{\mathbf{x},\mathcal{D}_{N}})$. This can be accomplished using class probability estimation (CPE) on the labels $z:=\mathds{1}[{y>\tau}]\in\{0,1\}$ so $p({y>\tau}|{\mathbf{x},\mathcal{D}_{N}})=p({z=1}|{\mathbf{x},\mathcal{D}_{N}})\approx\pi_{\theta}({\mathbf{x}})$, where $\pi_{\theta}:\mathcal{X}\to[0,1]$. We can estimate the class probabilities using a proper scoring rule (Gneiting & Raftery, 2007) such as log-loss,
$\mathcal{L_{\textrm{CPE}}}\!\left({\theta,\mathcal{D}_{N}^{z}}\right):=-\frac{1}{N}\sum\nolimits_{n=1}^{N}z_{n}\log\pi_{\theta}({\mathbf{x}_{n}})+(1-z_{n})\log(1-\pi_{\theta}({\mathbf{x}_{n}})),$ (9)
where $\mathcal{D}_{N}^{z}=\{(z_{n},\mathbf{x}_{n})\}^{N}_{n=1}$. The VSD objective and gradient estimator using CPE then become,
@@
$\nabla_{\phi}\mathcal{L_{\textrm{ELBO}}}\!\left({\phi,\theta}\right)$ $=\mathbb{E}_{q({\mathbf{x}}|{\phi})}\!\left[{\left(\log\pi_{\theta}({\mathbf{x}})-\log q({\mathbf{x}}|{\phi})+\log p({\mathbf{x}}|{\mathcal{D}_{0}})\right)\nabla_{\phi}\log q({\mathbf{x}}|{\phi})}\right].$ (11)
into which we plug $\theta_{t}^{*}=\operatorname*{arg\,\!min}_{\theta}\mathcal{L_{\textrm{CPE}}}\!\left({\theta,\mathcal{D}_{N}^{z}}\right)$. We refer to this strategy as CPE-PI. Using a CPE enables the use of more scalable estimators than GP-PI, satisfying our desideratum (D3) Scalable. This is crucial if we choose to run more than a few rounds of experiments with $B={\mathcal{{O}}}(1000)$. Since VSD is a black-box method we may choose to use CPEs that are non-differentiable, such as decision tree ensembles. The complete VSD algorithm is given in Algorithm 1 and depicted in Figure B.1. We have allowed for a threshold function, $\tau_{t}=f_{\tau}\!\left({\{y:y\in\mathcal{D}_{N}\},\gamma_{t}}\right)$, that can be used to modify the threshold each round. For example, an empirical quantile function $\tau_{t}=\hat{Q}_{y}(\gamma_{t})$ where $\gamma_{t}\in(0,1)$ as in Tiao et al. (2021). Or a constant $\tau$ for estimating a constant distribution of the super level-set.
Algorithm 1 VSD optimization loop with CPE-PI.
1: Threshold function $f_{\tau}$ and $\gamma_{1}$, dataset $\mathcal{D}_{N}$, black-box ${f\!\centerdot}$, prior $p({\mathbf{x}}|{\mathcal{D}_{0}})$, CPE $\pi_{\theta}({\mathbf{x}})$, variational family $q({\mathbf{x}}|{\phi})$, budget $T$ and $B$.
@@
3:     $\mathcal{D}^{z}_{N}$leftarrow $\{(z_{n},\mathbf{x}_{n})\}^{N}_{n=1}$,  where $z_{n}=\mathds{1}[{y_{n}>\tau}]$
@@
12:     $\mathcal{D}_{N}$leftarrow $\mathcal{D}_{N}\cup\{(\mathbf{x}_{bt},y_{bt})\}^{B}_{b=1}$